# Adversary Resilient Learned Bloom Filters

### v1

@@ 3 Secure Constructions for the Learned Bloom Filter
We attempt two solutions to the problem of securing the Learned Bloom Filter. Our first solution combines a Learned Bloom Filter technique called “sandwiching” introduced by Mitzenmacher [8] and the use of a pseudo-random permutation on the Bloom Filter input set first proposed by Naor and Yogev [3]. Our second solution relies on the (non Sandwiched) Learned Bloom Filter and instead employs the use of pseudo-random permutations to two Backup Classical Bloom Filters that we introduce as part of the construction.
In Section 3.1, we first provide a review of the setting from Section 2.2. Then we formally define a Sandwiched Learned Bloom Filter. Finally we provide a brief introduction to pseudo-random permutations. We introduce the Uptown Bodega Filter in Section 3.2. We then discuss concerns regarding the utility of the Uptown Bodega Filter which lead us to the Downtown Bodega Filter, which we introduce in Section 3.3.
@@ A Bloom Filter Learning Model $l_{r}^{\mathcal{D}}$ trained on $\mathcal{D}$
Proof.
The proof follows from the definition of a Sandwiched Learned Bloom Filter. ∎
We use the standard definitions for pseudo-random permutations. We provide a brief but self-contained treatment adapted from Chapter 3 of Katz Lindell [13] here. Let $\text{Perm}_{n}$ be the set of all permutations on $\{0,1\}^{n}$.
Definition 3.3.
Definition 3.4.
Definition 3.5.
Theorem $4.8$ of Naor Yogev [3] proves that for a classical steady $(n,\epsilon)$-Bloom Filter that uses $m$ bits of memory, if pseudo-random permutations exist, then there exists a negiligible function $\mathtt{negl}$ such that for security parameter $\lambda$ there exists a $(n,\epsilon+\mathtt{negl}(\lambda))$-adversarial resilient Bloom Filter that uses $m^{{}^{\prime}}=m+\lambda$ bits of memory. This secure Bloom Filter can be constructed by running the initialization algorithm on $S^{{}^{\prime}}={\{F_{k}(x):x\in S\}}$ instead of $S$ [3].
3.2 The Uptown Bodega Filter
Now, we introduce our first secure construction for the Learned Bloom Filter built on top of the Sandwiched Learned Bloom Filter and a pseudo-random permutation. We call this construction an Uptown Bodega Filter.
Definition 3.6 (Uptown).
Theorem 3.7.
Our proof follows the same line of reasoning as Theorem $4.8$ of Naor Yogev [3] which proves that, if pseudo-random permutations exist, any probabilistic polynomial time adversary cannot distinguish between the Classical Bloom Filter with pseudo-random permutations and Challenge 1 by more than a negligible advantage. Due to Lemma 3.2, this result also holds for the Uptown Bodega Filter. The crux of the proof relies on the fact that running $F_{k}$ on an adversary’s queries permits us to consider the queries as random and not chosen adaptively by an adversary, while having no effect on the correctness of the Uptown Bodega Filter.
Proof.
We construct a Uptown Bodega Filter $UB_{k,r}$ from a Sandwiched Bloom Filter $SB_{r}$ as follows. Choose the first key $k\in{\{0,1\}}^{\lambda}$ from pseudo-random permutation, $F_{k}$, over ${\{0,1\}}^{\log|U|}$. Let $S_{k}=F_{k}(S)={\{F_{k}(x):x\in S\}}$. We initialize $SB_{R}$ with $S_{k}$. For each input $x$, our query algorithm outputs $SB_{k,r}(F_{k}(x))$. Just like in the classical Bloom Filter construction of [3], the only additional memory required is to store the key $k$ for $F_{k}$, which is $\lambda$ bits. Similarly, the running time of the query algorithm of the Uptown Bodega Filter is one pseudo-random permutation more than the running time of the query algorithm of the Sandwiched Learned Bloom Filter.
Now we show that no probabilistic polynomial time (PPT) adversary $A$ can distinguish between the Uptown Bodega Filter we constructed using $R(\cdot)$ and the Uptown Bodega Filter construction that uses the pseudo-random permutation $F_{k}$ by more than a negligible advantage. Suppose that there does exist a non-negligible function $\delta(\lambda)$ such that $A$ can attack $UB_{k,r}$ and find a false positive with probability $\epsilon+\delta(\lambda)$. We can run $A$ on $UB_{k,r}$ where the oracle is replaced by an oracle that is either random or pseudo-random. We return $1$ if $A$ successfully finds a false positive. This implies that we may distinguish between a truly random permutation and a pseudo-random permutation with probability $\geq\delta(\lambda)$. This contradicts the indistinguishability of pseudo-random permutations. ∎
3.2.1 Utility of the Uptown Bodega Filter
While we have shown the existence of $(n,\epsilon)$-adversarial resilient Uptown Bodega Filters, it is unclear whether there is any utility in using an Uptown Bodega Filter over a secure Classical Bloom Filter. This is due to the fact that we train our learning model on $F_{k}(x)\in S$, a pseudo-random permutation of the set $S$, instead of training it directly on $S$. Intuitively, any structure in $S$ that can be “learned” by a Bloom Filter Learning Model is destroyed by taking the pseudo-random permutation of $S$. In other words, if pseudo-random permutations exist, then to any Bloom Filter Learning Model that runs in Probabilistic Polynomial Time (PPT), $F_{k}(x)\in S$ should appear the same as $|S|$ values sampled from a truly random oracle $R(\cdot)$. A precise proof for this is beyond the scope of this paper.
@@ 3.3 The Downtown Bodega Filter
To address the utility concerns of the Uptown Bodega Filter, we introduce a second Secure Learned Bloom Filter which we call the Downtown Bodega Filter.
Figure 4: A Downtown Bodega Filter. Both Positives and Negatives results for the Learned Oracle are routed to Backup Bloom Filters secured with Pseudorandom Permutations
@@ A Bloom Filter Learning Model $l_{r}^{\mathcal{D}}$ trained on $\mathcal{D}$
Lemma 3.9 is just a rephrasing of Theorem $4.8$ of Naor Yogev [3] and follows directly from the theorem. We include a self-contained proof here for completeness. We also make it more evident that the proof holds not just for adversarial resilient Bloom Filters but also for adversarial reveal resilient Bloom Filters that provide guarantees under a strictly stronger adversarial model. It is important to note that, while the adversary has access to the internal representation of the Downtown Bodega Filter, the adversary does not have access to the secret keys $k_{A}$ and $k_{B}$. This is consistent with the formulation of Naor Yogev [3] who state at the end of Section 4 in their paper, with reference to Theorem $4.8$, “Notice that, in all the above constructions only the pseudo-random function (permutation) key must remain secret. That is, we get the same security even when the adversary gets the entire memory of the Bloom filter except for the PRF (PRP) key.”
Proof.
@@
Now we show that no probabilistic polynomial time (PPT) adversary $A$ can distinguish between the Backup Bloom Filter $A$ we constructed using $\mathcal{R}(\cdot)$ and the Backup Classical Bloom Filter $A$ construction that uses the pseudo-random permutation $F_{k_{A}}$ by more than a negligible advantage. Suppose that there does exist a non-negligible function $\delta(\lambda)$ such that adversary $A$ can attack Backup Classical Bloom Filter $A$ and find a false positive with probability $\epsilon+\delta(\lambda)$. We can run adversary $A$ on a Backup Classical Bloom Filter $A$ where the oracle is replaced by an oracle that is either random or pseudo-random. We return $1$ if $A$ successfully finds a false positive. This implies that we may distinguish between a truly random permutation and a pseudo-random permutation with probability $\geq\delta(\lambda)$. This contradicts the indistinguishability of pseudo-random permutations. ∎
@@
Therefore, for any probabilistic polynomial time (PPT) adversary to induce a false positive in the overall Downtown Bodega Filter construction, they must either induce a false positive in Backup Classical Bloom Filter $A$ or Backup Classical Bloom Filter $B$. We have already proven in Lemma 3.9 that both Backup Classical Bloom Filter $A$ and Backup Classical Bloom Filter $B$ are $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient. It follows that the entire construction is $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient. This concludes our proof. ∎
@@ 4 Discussion
In this section, we discuss the only two known attacks on the Learned Bloom Filter, introduced by Reviriego et al [2]. We refer to Attack 1 from their work as the Blackbox Mutation Attack, and Attack 2 from their work as the Whitebox Mutation Attack respectively. We discuss the Blackbox Mutation Attack in Section 4.1, and the Whitebox Mutation Attack in Section 4.2. Both sections include details on how our Secure Learned Bloom Filter constructions mitigate the attack.
@@ 4.1 Black-box Mutation Attack
The black-box adversarial model defined by Reviriego et al [2] is slightly weaker but very similar to the adversarial model we define in Challenge 3 of Section 2.2. Both our adversarial model and Reviriego et al’s black-box adversary model allows the adversary access to query the Learned Bloom Filter. One major difference is that our model allows the adversary to choose the initial set $S$ that is represented by the Learned Bloom Filter, whereas Reviriego et al’s model does not. In their attack, Reviriego et al first test elements until a positive (whether a false positive or true positive) is found. They then mutate the positive by changing a small fraction of the bits in the input in order to generate more false positives. The attack targets the Learned Oracle (recall Figure 2) by making it generate false positives without the input reaching the Backup Bloom Filter.
Since, as we mentioned, Reviriego et al’s attack relies on the fact that the adversary can query the Learned Oracle and make it generate false positives without any involvement of the Backup Bloom Filter, a Sandwiched Bloom Filter construction is intuitively more resilient against this kind of mutation attack. Since the Uptown Bodega Filter takes a pseudo-random permutation of the input $x$ prior to sending it to either the Learned Oracle or the Bloom Filters, the mutation approach is no longer viable. Unlike in the Uptown Bodega Filter, the Downtown Bodega Filter does give the adversary the ability to query the Learned Oracle directly. However, the extra (pseudo-random permuted) Backup Classical Bloom Filter for the case $l(x)>\tau$ lowers the probability of mutation-based false positives from the Learned Oracle ending up as false positives in the overall construction.
@@ 4.2 White-box Mutation Attack
The white-box adversarial model defined by Reviriego et al [2] is similar to the adversarial model we define in Challenge 4 in Section 2.2. With knowledge of the state of the Learned Oracle, the adversary can generate mutations in a more sophisticated way. Reviriego et al provide the example of a malicious URL dataset where an adversary may begin with a non-malicious URL and make changes such as removing the “s” in “https” or removing the “www” to generate false positives. Since we have shown the Downtown Bodega Filter to be $(n,t,\epsilon)$-adversarial reveal resilient in the steady setting, such mutations will not provide the adversary any advantage over the construction.
@@ 5 Hybrid Adversarial Model
In this section, we first define a hybrid model, where part of the queries are chosen by an adversary, while the rest are non-adversarial (“regular”) queries. Next we analyze the performance of the Downtown Bodega Filter and the Secure Classical Bloom Filter respectively in our hybrid model. We provide results for the conditions in which a Downtown Bodega Filter outperforms the Secure Classical Bloom Filter in the hybrid model. We then discuss the tradeoffs of our approach and provide a realistic example where the Downtown Bodega Filter displays better performance as compared to a Secure Classical Bloom Filter under many settings.
@@ 5.1 Hybrid Model
Let $A=(A_{C},A_{Q})$ be a probabilistic polynomial time (PPT) adversary as defined in Section 2. Consider a set of $N$ queries sent to a Bloom Filter. Our adversary, $A$, is allowed to choose exactly $\alpha N$ of those queries, where $\alpha\in[0,1]$. As an exampleto, the $N$ queries may be part of a streaming workload under any of the streaming models described by Muthukrishnan [14].
@@ 5.2 Downtown Bodega Filter
Figure 5: To generate a false positive in the Downtown Bodega Filter, the adversary must either 1) generate a false positive in the Learned Oracle and direct their query through Backup Bloom Filter $A$ or 2) generate a true negative in the Learned Oracle and direct their query through Backup Bloom Filter $B$
@@
    The query generates a false positive in the Learned Oracle and a false positive in Backup Bloom Filter $A$.
@@
    The query generates a true negative in the Learned Oracle and a false positive in Backup Bloom Filter $B$
We first model the false positive probability of the Downtown Bodega Filter as a function of its memory budget, $m$, the set that the Downtown Bodega Filter is encoding, $S$, and the corresponding training dataset $\mathcal{D}$. Without loss of generality, let $\alpha_{P}$ of the adversarial queries be such that they generate false positives in the Learned Oracle and go through Backup Bloom Filter $A$. Similarly, let $\alpha_{N}$ of the adversarial queries be such that they generate true negatives and go through Backup Bloom Filter $B$. Note that $\alpha=\alpha_{P}+\alpha_{N}$.
Let $\text{FPR}(S,m)$ be the expected false positive probability of a Classical Bloom Filter that encodes the set $S$ with memory budget $m$. Let $\text{FPR}_{L}(S,\mathcal{D},m)$ be the expected false positive probability of a Learned Oracle $L$ that encodes the set $S$ using the training dataset $\mathcal{D}$ with memory budget $m$. Similarly, let $\text{TNR}_{L}(S,\mathcal{D},m)$ be the expected true negative probability of a Learned Oracle $L$ that encodes the set $S$ using the training dataset $\mathcal{D}$ with memory budget $m$. We make the assumption that the correctness probability of the Learned Oracle is independent of the correctness probability of the Backup Classical Bloom Filters. In particular, we assume that for any $m,m^{\prime}$, $\text{FPR}_{L}(m)\cap\text{FPR}(m^{\prime})=\text{FPR}_{L}(m)\text{FPR}(m^{\prime})$ and $\text{TNR}_{L}(m)\cap\text{FPR}(m^{\prime})=\text{TNR}_{L}(m)\text{FPR}(m^{\prime})$.
Consider a system where the total memory budget is $M$. Let the memory allocation of a Downtown Bodega Filter from memory budget $M$ be assigned as follows. Let $m_{L}$ be the number of bits of memory assigned to a Learned Oracle $L$. Let $m_{A}$ be the number of bits of memory assigned to Backup Bloom Filter $A$. Let $m_{B}$ be the number of bits of memory assigned to Backup Bloom Filter $B$. Let $\lambda$ be the number of bits assigned to the key of the pseudo-random permutations used prior to Backup Bloom Filter $A$ and Backup Bloom Filter $B$. Note that to stay within the memory budget it must hold that $M\geq m_{L}+m_{A}+m_{B}+2\lambda$.
@@
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns positive, and where the probability is taken over the random coins of the pseudo-random permutations, the random coins used in the construction of the Learned Oracle, the random coins used in the construction of Backup Classical Bloom Filters $A$ and $B$, and the random coins used in the generation of the non-adversarial queries.
Proof.
From the definition of a Downtown Bodega Filter (Definition 3.8, in particular look at the formulation for the query algorithm $\tilde{Q}_{s}$. Figure 5 is also helpful here) it follows that a false positive in the overall construction must either be a false positive in Backup Classical Bloom Filter $A$ or a false positive in Backup Classical Bloom Filter $B$. If the query is a false positive in Backup Classical Bloom Filter $A$, it must also be a false positive in the Learned Oracle. Alternatively, if the query is a false positive in Backup Classical Bloom Filter $B$, it must be a true negative in the Learned Oracle. The result follows. ∎
We now derive an expression for the false positive probability of the Downtown Bodega Filter in the hybrid adversarial setting. Recall that we are assuming that out of $N$ queries, the adversary makes $\alpha_{P}$ queries that generate a false positive in the Learned Oracle and $\alpha_{N}$ queries that generate a true negative in the Learned Oracle.
@@
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Learned Oracle $L$ returns positive, and where the probability is taken over the random coins of the pseudo-random permutations, the random coins used in the construction of the Learned Oracle, the random coins used in the construction of Backup Classical Bloom Filters $A$ and $B$, and the random coins used in the generation of the non-adversarial queries.
Proof.
@@
Case 2: The query is an adversary generated query such that it generates a false positive in the Learned Oracle. Since the Learned Oracle generating a false positive and the Learned Oracle generating a true negative are mutually exclusive events, the False Positive Probability for the query is the False Positive Probability of Backup Classical Bloom Filter $A$ i.e $\text{FPR}(S^{\prime\prime},m_{A})$. There are $\alpha_{P}N$ such queries.
Case 3: The query is an adversary generated query such that it generates a true negative in the Learned Oracle. Following logic similar to case 2, we can derive the False Positive Probability of the query to be $\text{FPR}(S^{\prime},m_{B})$. There are $\alpha_{N}N$ such queries.
The expected false positive probability of $N$ queries is then $\frac{1}{N}(\alpha_{P}N\cdot\text{FPR}(S^{\prime\prime},m_{A})+\alpha_{N}N\cdot\text{FPR}(S^{\prime\prime},m_{B})+(1-\alpha_{P}-\alpha_{N})N\cdot\text{FPR}_{DB}(S,M))$. The statement of the theorem follows. ∎

### v3

@@ 3 Secure Constructions for the Learned Bloom Filter
In this section, we propose a solution to the problem of securing the Learned Bloom Filter. We call our construction the Downtown Bodega Filter. Section 3.1 contains the background necessary to construct the Downtown Bodega Filter. We formally define the Downtown Bodega Filter in Section 3.2. The interested reader may also refer to Appendix A which discusses an alternate construction of a secure Learned Bloom Filter called the Uptown Bodega Filter.
@@ 3.1 Preliminaries
We introduce a Learned Bloom Filter technique called “sandwiching” introduced by Mitzenmacher [Mit18]. We then discuss the use of a pseudo-random permutation on the Bloom Filter input set first proposed by Naor Yogev [NE19].
@@ A Bloom Filter Learning Model $l_{r}^{\mathcal{D}}$ trained on $\mathcal{D}$
Proof 3.3.
The proof follows from the definition of a Sandwiched Learned Bloom Filter.
We use the standard definitions for pseudo-random permutations in this work. We provide a brief but self-contained treatment of pseudo-random permutations adapted from Chapter 3 of Katz Lindell [KL14] in Appendix B. Theorem $4.8$ of Naor Yogev [NE19] proves that for a classical steady $(n,\epsilon)$-Bloom Filter that uses $m$ bits of memory, if pseudo-random permutations exist, then there exists a negiligible function $\mathtt{negl}$ such that for security parameter $\lambda$ there exists a $(n,\epsilon+\mathtt{negl}(\lambda))$-adversarial resilient Classical Bloom Filter that uses $m^{{}^{\prime}}=m+\lambda$ bits of memory. This secure Classical Bloom Filter can be constructed by running the initialization algorithm on $S^{{}^{\prime}}={\{F_{k}(x):x\in S\}}$ instead of $S$ [NE19].
@@ 3.2 The Downtown Bodega Filter
We introduce a secure construction of the Learned Bloom Filter which we call the Downtown Bodega Filter.
Figure 4: A Downtown Bodega Filter. Both Positives and Negatives results for the Bloom Filter Learning Model are routed to Backup Bloom Filters secured with Pseudo-random Permutations
@@ A Bloom Filter Learning Model $l_{r}^{\mathcal{D}}$ trained on $\mathcal{D}$
Lemma 3.5 is just a rephrasing of Theorem $4.8$ of Naor Yogev [NE19] and follows directly from the theorem. We include a self-contained proof here for completeness. We also make it more evident that the proof holds not just for adversarial resilient Bloom Filters but also for adversarial reveal resilient Bloom Filters that provide guarantees under a strictly stronger adversarial model. It is important to note that, while the adversary has access to the internal representation of the Downtown Bodega Filter, the adversary does not have access to the secret keys $k_{A}$ and $k_{B}$. This is consistent with the formulation of Naor Yogev [NE19] who state at the end of Section 4 in their paper, with reference to Theorem $4.8$, “Notice that, in all the above constructions only the pseudo-random function (permutation) key must remain secret. That is, we get the same security even when the adversary gets the entire memory of the Bloom filter except for the PRF (PRP) key.”
Proof 3.6.
@@
Now we show that no probabilistic polynomial time (PPT) adversary $A$ can distinguish between the Backup Bloom Filter $A$ we constructed using $\mathcal{R}(\cdot)$ and the Backup Classical Bloom Filter $A$ construction that uses the pseudo-random permutation $F_{k_{A}}$ by more than a negligible advantage. Suppose that there does exist a non-negligible function $\delta(\lambda)$ such that adversary $A$ can attack Backup Classical Bloom Filter $A$ and find a false positive with probability $\epsilon+\delta(\lambda)$. We can run adversary $A$ on a Backup Classical Bloom Filter $A$ where the oracle is replaced by an oracle that is either random or pseudo-random. We return $1$ if $A$ successfully finds a false positive. This implies that we may distinguish between a truly random permutation and a pseudo-random permutation with probability $\geq\delta(\lambda)$. This contradicts the indistinguishability of pseudo-random permutations.
@@
Therefore, for any probabilistic polynomial time (PPT) adversary to induce a false positive in the overall Downtown Bodega Filter construction, they must either induce a false positive in Backup Classical Bloom Filter $A$ or Backup Classical Bloom Filter $B$. We have already proven in Lemma 3.5 that both Backup Classical Bloom Filter $A$ and Backup Classical Bloom Filter $B$ are $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient. It follows that the entire construction is $(n,\epsilon+\mathrm{negl}(\lambda))$-adversarial reveal resilient. This concludes our proof.
@@ 4 Discussion
In this section, we discuss the only two known attacks on the Learned Bloom Filter, introduced by Reviriego et al [RAHDS21]. We refer to Attack 1 from their work as the Blackbox Mutation Attack, and Attack 2 from their work as the Whitebox Mutation Attack respectively. We discuss the Blackbox Mutation Attack in Section 4.1, and the Whitebox Mutation Attack in Section 4.2. Both sections include details on how our Secure Learned Bloom Filter construction mitigates the attack.
@@ 4.1 Black-box Mutation Attack
The black-box adversarial model defined by Reviriego et al [RAHDS21] is slightly weaker but very similar to the adversarial model we define in Challenge 3 of Section 2.2. Both our adversarial model and Reviriego et al’s black-box adversary model allows the adversary access to query the Learned Bloom Filter. One major difference is that our model allows the adversary to choose the initial set $S$ that is represented by the Learned Bloom Filter, whereas Reviriego et al’s model does not. In their attack, Reviriego et al first test elements until a positive (whether a false positive or true positive) is found. They then mutate the positive by changing a small fraction of the bits in the input in order to generate more false positives. The attack targets the Bloom Filter Learning Model (recall Figure 2) by making it generate false positives without the input reaching the Backup Bloom Filter. This attack is mitigated by the Downtown Bodega Filter because the Downtown Bodega Filter passes queries through a Secure Classical Backup Bloom Filter even for the case where the Bloom Filter Learning Model returns true i.e when $l(x)>\tau$.
@@ 4.2 White-box Mutation Attack
The white-box adversarial model defined by Reviriego et al [RAHDS21] is similar to the adversarial model we define in Challenge 4 in Section 2.2. With knowledge of the state of the Bloom Filter Learning Model, the adversary can generate mutations in a more sophisticated way. Reviriego et al provide the example of a malicious URL dataset where an adversary may begin with a non-malicious URL and make changes such as removing the “s” in “https” or removing the “www” to generate false positives. Since we have shown the Downtown Bodega Filter to be $(n,t,\epsilon)$-adversarial reveal resilient in the steady setting, such mutations will not provide the adversary any advantage over the construction.
@@ 5 Hybrid Adversarial Model
In this section, we first define a hybrid model, where part of the queries are chosen by an adversary, while the rest are non-adversarial (“regular”) queries. Next we analyze the performance of the Downtown Bodega Filter and the Secure Classical Bloom Filter respectively in our hybrid model. We provide results for the conditions in which the Downtown Bodega Filter outperforms the Secure Classical Bloom Filter in the hybrid model. We then discuss the trade-offs of our approach and provide a realistic example where the Downtown Bodega Filter displays better performance as compared to the Secure Classical Bloom Filter under many settings.
@@ 5.1 Hybrid Model
Let $A=(A_{C},A_{Q})$ be a probabilistic polynomial time (PPT) adversary as defined in Section 2. Consider a set of $N$ queries sent to a (Classical or Learned) Bloom Filter. Our adversary, $A$, is allowed to choose exactly $\alpha N$ of those queries, where $\alpha\in[0,1]$. As an example, the $N$ queries may be part of a streaming workload under any of the streaming models described by Muthukrishnan [Mut05].
@@ 5.2 Downtown Bodega Filter
Figure 5: To generate a false positive in the Downtown Bodega Filter, the adversary must either 1) generate a false positive in the Bloom Filter Learning Model and direct their query through Backup Classical Bloom Filter $A$ or 2) generate a true negative in the Bloom Filter Learning Model and direct their query through Backup Classical Bloom Filter $B$
@@
    The query generates a false positive in the Bloom Filter Learning Model and a false positive in Backup Bloom Filter $A$.
@@
    The query generates a true negative in the Bloom Filter Learning Model and a false positive in Backup Bloom Filter $B$
We first model the false positive probability of the Downtown Bodega Filter as a function of its memory budget, $m$, the set that the Downtown Bodega Filter is encoding, $S$, and the corresponding training dataset $\mathcal{D}$. Without loss of generality, let $\alpha_{P}$ of the adversarial queries be such that they generate false positives in the Bloom Filter Learning Model and go through Backup Bloom Filter $A$. Similarly, let $\alpha_{N}$ of the adversarial queries be such that they generate true negatives and go through Backup Bloom Filter $B$. Note that $\alpha=\alpha_{P}+\alpha_{N}$.
Let $\text{FPR}(S,m)$ be the expected false positive probability of a Classical Bloom Filter that encodes the set $S$ with memory budget $m$. Let $\text{FPR}_{L}(S,\mathcal{D},m)$ be the expected false positive probability of a Bloom Filter Learning Model $L$ that encodes the set $S$ using the training dataset $\mathcal{D}$ with memory budget $m$. Similarly, let $\text{TNR}_{L}(S,\mathcal{D},m)$ be the expected true negative probability of a Bloom Filter Learning Model $L$ that encodes the set $S$ using the training dataset $\mathcal{D}$ with memory budget $m$. We make the assumption that the correctness probability of the Bloom Filter Learning Model is independent of the correctness probability of the Backup Classical Bloom Filters. In particular, we assume that for any $m,m^{\prime}$, $\text{FPR}_{L}(m)\cap\text{FPR}(m^{\prime})=\text{FPR}_{L}(m)\text{FPR}(m^{\prime})$ and $\text{TNR}_{L}(m)\cap\text{FPR}(m^{\prime})=\text{TNR}_{L}(m)\text{FPR}(m^{\prime})$.
Consider a system where the total memory budget is $M$. Let the memory allocation of a Downtown Bodega Filter from memory budget $M$ be assigned as follows. Let $m_{L}$ be the number of bits of memory assigned to a Bloom Filter Learning Model $L$. Let $m_{A}$ be the number of bits of memory assigned to Backup Bloom Filter $A$. Let $m_{B}$ be the number of bits of memory assigned to Backup Bloom Filter $B$. Let $\lambda$ be the number of bits assigned to the key of the pseudo-random permutations used prior to Backup Bloom Filter $A$ and Backup Bloom Filter $B$. Note that to stay within the memory budget it must hold that $M\geq m_{L}+m_{A}+m_{B}+2\lambda$.
@@
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Bloom Filter Learning Model $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Bloom Filter Learning Model $L$ returns positive, and where the probability is taken over the random coins of the pseudo-random permutations, the random coins used in the construction of the Bloom Filter Learning Model, the random coins used in the construction of Backup Classical Bloom Filters $A$ and $B$, and the random coins used in the generation of the non-adversarial queries.
Proof 5.2.
From the definition of a Downtown Bodega Filter (Definition 3.4, in particular look at the formulation for the query algorithm $\tilde{Q}_{s}$. Figure 5 is also helpful here) it follows that a false positive in the overall construction must either be a false positive in Backup Classical Bloom Filter $A$ or a false positive in Backup Classical Bloom Filter $B$. If the query is a false positive in Backup Classical Bloom Filter $A$, it must also be a false positive in the Bloom Filter Learning Model. Alternatively, if the query is a false positive in Backup Classical Bloom Filter $B$, it must be a true negative in the Bloom Filter Learning Model. The result follows.
We now derive an expression for the false positive probability of the Downtown Bodega Filter in the hybrid adversarial setting. Recall that we are assuming that out of $N$ queries, the adversary makes $\alpha_{P}$ queries that generate a false positive in the Bloom Filter Learning Model and $\alpha_{N}$ queries that generate a true negative in the Bloom Filter Learning Model.
@@
where $S^{\prime}\subset S$ is the set of elements in $S$ for which the Bloom Filter Learning Model $L$ returns negative, $S^{\prime\prime}\subset S$ is the set of elements in $S$ for which the Bloom Filter Learning Model $L$ returns positive, and where the probability is taken over the random coins of the pseudo-random permutations, the random coins used in the construction of the Bloom Filter Learning Model, the random coins used in the construction of Backup Classical Bloom Filters $A$ and $B$, and the random coins used in the generation of the non-adversarial queries.
Proof 5.4.
@@
Case 2: The query is an adversary generated query such that it generates a false positive in the Bloom Filter Learning Model. Since the Bloom Filter Learning Model generating a false positive and the Bloom Filter Learning Model generating a true negative are mutually exclusive events, the False Positive Probability for the query is the False Positive Probability of Backup Classical Bloom Filter $A$ i.e $\text{FPR}(S^{\prime\prime},m_{A})$. There are $\alpha_{P}N$ such queries.
Case 3: The query is an adversary generated query such that it generates a true negative in the Bloom Filter Learning Model. Following logic similar to case 2, we can derive the False Positive Probability of the query to be $\text{FPR}(S^{\prime},m_{B})$. There are $\alpha_{N}N$ such queries.
The expected false positive probability of $N$ queries is then $\frac{1}{N}(\alpha_{P}N\cdot\text{FPR}(S^{\prime\prime},m_{A})+\alpha_{N}N\cdot\text{FPR}(S^{\prime\prime},m_{B})+(1-\alpha_{P}-\alpha_{N})N\cdot\text{FPR}_{DB}(S,M))$. The statement of the theorem follows.